Allison:

The radius of a circle is 10 ft. Find its area in terms of π.

Joe348:

what are the first 3 numbers of pi?

Allison:

3.14...?

Joe348:

yes

AZ:

that doesn't even help her at all smh

Joe348:

im trying too tho

AZ:

It's really simple \( \text{Area} = \pi r^2\) so just replace r with 10 what is 10^2

Allison:

20

Allison:

40

AZ:

test why does enters not work why is it all one line??? and no 10^2 is 10 * 10

Allison:

Huh?

Allison:

100

AZ:

so essentially \( area = 10^2 \pi = \boxed{100\pi}\)

AZ:

that's your final answer, you don't use 3.14 for pi. You just leave it in terms of pi

AZ:

Does that make sense?

Joe348:

but area= pi times 10ft ^2 so basically it is pi without the decimal

AZ:

lol where did decimals even come from

Allison:

Pi without the decimal is just a big number.. 3141592..